A toy glider is tossed into the air. Its height, h, in metres, after t seconds is modelled by h = 9t – 3t2. Find the maximum height the toy glider reaches. Show the steps of your solution.

Answer by nerdybill(7384)

A toy glider is tossed into the air. Its height, h, in metres, after t seconds is modelled by h = 9t – 3t2. Find the maximum height the toy glider reaches. Show the steps of your solution.
.
The equation:
h = 9t – 3t^2
can be written as:
h = –3t^2 + 9t
is a parabola that opens downwards (based on the negative coefficient associated with the t^2 term). This means that the max will be at the vertex:
Find t value of the vertex:
t = -b/(2a)
t = -9/(2(-3))
t = -9/(-6)
t = 3/2 seconds
.
To find the height, we plug the value above back into the original equation:
h = –3t^2 + 9t
h = –3(3/2)^2 + 9(3/2)
h = –3(9/4) + 9(3/2)
h = –27/4 + 27/2
h = –27/4 + 54/4
h = (–27+54)/4
h = 27/4
h = 6.75 meters